OpenStudy (luigi0210):

Limit question:

OpenStudy (luigi0210):

\[\LARGE \lim_{x \rightarrow -\infty}(\sqrt{x^2+x+1}+x) \]

OpenStudy (dls):

Rationalize.

OpenStudy (dan815):

-inf

OpenStudy (dls):

-1/2 ?

OpenStudy (dan815):

my guess was wrong :O??

OpenStudy (dls):

no doubt

OpenStudy (dan815):

lol

OpenStudy (luigi0210):

\[\LARGE \lim_{x \rightarrow -\infty}(\sqrt{x^2+x+1}+x)*(\sqrt{x^2+x+1}-x)\] \[\LARGE \lim_{x \rightarrow -\infty} (x^2+x-1-x^2)\] \[\LARGE \lim_{x \rightarrow -\infty}(x-1)\] Did I screw up? xD

OpenStudy (luigi0210):

Whoops *+1

OpenStudy (dls):

did you forget how to rationalize ? o.O you just multiplied but didn't divide.

OpenStudy (dan815):

it's -1/2

OpenStudy (luigi0210):

Yes I did xD \[\LARGE \lim_{x \rightarrow -\infty} \frac{(x+1)}{\sqrt{x^2+x+1}-x}\]

OpenStudy (dan815):

okay now divide top and bottom by x

OpenStudy (dls):

like he said

OpenStudy (dls):

for denominator i'd help a bit \[\LARGE \sqrt{(x^2(1+\frac{1}{x}+\frac{1}{x^2})} -x \]

OpenStudy (dls):

x^2 will become x outside and then x will cancel out Then use the fact that............ \[\Huge \lim_{x \rightarrow -\inf} \frac{1}{x^2} ,\frac{1}{x} ->0\]

OpenStudy (dan815):

=[ how do i type in latex

OpenStudy (luigi0210):

You forgot your "\"

OpenStudy (dan815):

\[hey\]

OpenStudy (luigi0210):

Alright, thank you guys :)

OpenStudy (dls):

uw :)